# Targeted Nakamoto: A Bitcoin Protocol to Balance Network Security and Energy Consumption

### v3

@@ Abstract
In a Proof-of-Work blockchain such as Bitcoin mining hashrate is increasing in the block reward. An increase in hashrate reduces network vulnerability to attack (a reduction in security cost) while increasing carbon emissions and electricity cost (an increase in externalities cost). This implies a tradeoff in total cost at different levels of hashrate and the existence of a hashrate interval where total cost is minimized. Targeted Nakamoto is a Proof-of-Work protocol augmentation that incentivizes miners to home in on a target hashrate interval. When hashrate is above target a ceiling is placed on the block reward a miner can receive. When hashrate is below target a floor is placed underneath the miner’s block reward. Monetary neutrality is maintained by a proportional increase in spending potential among addresses holding UTXO’s to match a deduction from total block reward when the ceiling is operative and a proportional reduction in spending potential among addresses holding UTXO’s to match an increase over the total block reward when the floor is binding.
Bitcoin faces two, diametrically opposed and temporally divergent, existential risks. Today, the massive energy consumed by Bitcoin mining is a focus of a political backlash aimed at de-legitimizing Bitcoin and restricting its use. In the future, as the minting of new UTXO’s to reward miners decreases toward zero, the revenue from mining will likely shrink, causing miners to exit and energy consumption to contract. The security of the network may be compromised by the low cost of an attack. In effect, Bitcoin’s energy consumption will be too low. Targeted Nakamoto is a modification to the Bitcoin protocol that is designed to avert these twin potential disasters.
@@ 1 Introduction
Proof-of -Work (PoW) blockchain cryptocurrencies like Bitcoin require the application of computing power by miners to operate the network. Miners assemble blocks and compete to solve a puzzle set by the code. The number of puzzle guesses (one hash per guess) a mining computer makes in a specified interval of time is its hashrate, which consumes electricity. Miners pay the direct cost of electricity. There are, in addition, two externalities created by mining activity. One externality is the vulnerability of the blockchain network to attack. A measure of vulnerability is the cost of forking the blockchain, which is related to the hashrate of active miners. The higher is hashrate, the higher is the cost of forking a PoW blockchain, which implies that network security is increasing in hashrate (equivalently, security cost is decreasing in hashrate). The other externality arises from the electricity consumed by hashrate. Miner demand for electricity increases both the price and aggregate consumption of electricity. The former is a cost imposed on competing users of electricity. The latter generates carbon emissions, which has a social cost. These electricity related costs are increasing in hashrate. The interaction of security risk and electricity consumption implies a tradeoff in total cost, defined as the sum of the two externalities costs, at different levels of hashrate. An increase in hashrate reduces the vulnerability of the network (equivalently, security cost is decreased), while it increases carbon emissions and creates upward pressure on the price of electricity. There may be a point or interval of hashrate in which total cost is minimized. Current PoW protocols, which I refer to as ”Nakamoto”, do not restrict hashrate. An increase in the value of the block reward creates an incentive for miners to apply more hashrate. A decrease in the value of the block reward has the opposite effect. Equilibrium hashrate can range from zero to unbounded. I focus on Bitcoin in this paper, although the protocol applies to any PoW cryptocurrency. Figure 4 shows the tight correlation over time between percentage change in the USD value of the Bitcoin block reward and the percentage change in energy consumption from Bitcoin mining.
@@
Figure 1: Bitcoin Energy Consumption vs USD Mining Rewards. Source: CBECI33footnotemark: 3and Coinmetrics. [^4] https://ccaf.io/cbnsi/cbeci
44footnotetext: https://coinmetrics.io
@@
Targeted Nakamoto (hereafter ”Targeted”) modifies Nakamoto to create an incentive for miners to push hashrate toward an interval whenever it strays outside the interval. The miner’s minimum block reward is increased when hashrate is below the interval, which induces miners to apply more hashrate. The miner’s maximum block reward is decreased when hashrate is above the interval, which induces miners to apply less hashrate. Monetary neutrality is maintained by offsetting the increase and decrease in the miner’s block reward by opposite adjustments to the spending potential of addresses with UTXO value, in proportion to the value held by each address.
1.1 Two existential risks to Bitcoin
Targeted mitigates two latent, potentially existential threats to Bitcoin. One threat arises from the halving of the minting of UTXO’s (the ”coinbase”), which ensures that transaction fees will eventually become the primary source of block rewards. [^5] Historically, the trend USD/BTC exchange rate has risen sufficiently to offset the halving of the coinbase. However, this cannot continue indefinitely for two reasons. First, coinbase will eventually be zero. Second, maintaining a fixed USD value of the coinbase paid to miners as halving continues would require the inflation adjusted USD to BTC exchange rate to double every four years. This implies a growth rate that exceeds US (and world) GDP by orders of magnitude, which is dynamically unsustainable (Tirole [21]). A transactor is motivated to add a fee to incentivize a miner to include the transaction in its block. [^6] Huberman et.al. [12] A transactor has, at best, a muted incentive to include a fee for network security, which is a non-rival public good that benefits all participants. Historically, transaction fees have comprised less than 10% of total block rewards on average and there is no assurance that transaction fees will increase in the future. [^7] Data on USD block rewards and transaction fees for Bitcoin are compiled by Coinmetrics https://coinmetrics.io The scheduled decline in the coinbase, if not matched by an increase in transaction fees, will cause miners to reduce hashrate, which will lower the cost of an attack on the network. The other threat comes from the reality and the adverse political reaction to the fact that the electricity consumed by Bitcoin miners adds to carbon emissions. Targeted mitigates both of these risks by incentivizing miners to home in on a hashrate interval. This ensures that the cost of attacking the network will not become dangerously cheap while limiting the network’s contribution to carbon emissions and electricity cost. Finally, Targeted preserves aggregate monetary neutrality by proportionally adjusting the spending potential of Bitcoin addresses to offset the minting or burning of UTXO value that that arises from the provision of incentives to miners.
1.2 Links to code and web API
@@
Accompanying this paper is a link to a code-base for controlling Hashrate and a Web API with an empirically estimated hashrate function and a module that enables the user to choose a target hashrate interval and control parameters and to run counterfactual experiments with alternative time-paths of model variables. Both instantiations were developed by Kristian Praizner of MIT.
@@
Both are anonymized for this submission.
1.3 Limitation
The decrease in network security risk and the increase in electricity cost and carbon emissions as hashrate increases ensures, with a mild restriction on curvature, the existence of a point or interval at which the sum of the direct and externalities costs are minimized. [^8] The restriction is that the function defined as the sum of the costs is is concave. However, locating the minimum requires a numerical assignment of the costs of network security risk and carbon emissions, which lies outside the scope of this paper. A related issue is the tracking of changes in the relationship between hashrate and electricity cost. The efficiency of Bitcoin ASICS has improved over time, reducing the Kw per hash. Coinmetrics [5] estimates that the average electricity per hash among active Bitcoin ASIC’s decreased by half from 2018 to 2022. If improvements continue, a fixed target hashrate interval will be associated with lower electricity consumption and a lower cost of attack. The difficulty here is not the tracking of changes in per kw electricity cost and ASIC efficiency, but rather use of an oracle that it requires, which violates a value held by many in the Bitcoin community. Finally, the relation between electricity cost and security may be non-stationary. A change in the USD value of Bitcoin, or in the size distribution or other characteristics of transactions, may alter the security achieved at any level of hashrate. Against these limitations stand the existential threats mentioned above, which may convince the reader of the need to form consensus on a target interval notwithstanding the gaps in our knowledge.
@@ 1.5 The research question and design constraints
Targeted augments Nakamoto by creating an incentive for miners to home in on a hashrate interval. The augmentation comes close to adhering to formal constraints which reflect core values held by the Bitcoin community. [^11] What constitutes a core feature is subjective. I developed the list herein after consultation with a number of participants in the Bitcoin network, including developers, miners and exchange operators. Ultimately, the reader will judge for herself whether my list is complete. The constraints are partitioned into the consensus protocol layer (C constraints) and the transaction protocol layer (T- constraints).
@@ 1.6 Roadmap
The rest of the paper is organized as follows. Section 2 provides an overview of the two types of cost, the sum of which Targeted is designed to minimize; costs that increase with hashrate and those that decrease with hashrate. Section 3 casts Nakamoto and Targeted in a mechanism design framework in order to articulate the key features of each protocol and their interaction and to show how Targeted augments Nakamoto. Section 4 shows how the ratio of mining puzzle difficulty to blockchain growth rate is a sufficient statistic for hashrate and then derives the expression for equilibrium hashrate as a function of, inter alia, the value of the block reward. Section 5 uses the relationship between the puzzle difficulty/growth rate ratio and hashrate, along with hashrate equilibrium to construct the policy for adjusting the miner’s block reward to incentivize miners to push hashrate toward the target hashrate interval. Section 6 demonstrates that the hashrate control algorithm dynamically pushes hashrate toward the target interval. Section 7 states a protocol that renders the aggregate money supply and the distribution of spending potential unaffected by the Hashrate Control Algorithm. The paper concludes with a summary of the key properties of Targeted and call attention (again) to the urgency for the Bitcoin community to address the issues discussed in this paper.
2 Externalities Costs
This section explains the costs associated with honest miner hashrate. Section 2.1 reviews evidence of a link between Bitcoin hashrate and carbon emissions. Section 2.2 discusses three vectors of attack on the network that become more feasible as hashrate decreases; double-spend, eclipse and shorting attacks. In Section 2.3 derives a target hashrate interval as a function of externalities costs.
2.1 Externalities costs that increase with hashrate
Hashing, or puzzle guessing, involves computation that consumes electricity. This generates two external effects. One is to increase the price of electricity and the other is to increase the social cost of carbon emissions from the from fossil fuel burning electricity generators. [^12] Chen et.al. [14] model the flow of carbon emissions from multiple energy sources. These costs are related in the sense that an increase in hashrate will, by increasing electricity demand, put upward pressure on electricity price and, to the extent that the increase in demand is met by an increase in electricity generation (i.e. that electricity supply is not inelastic), cause a rise in carbon emissions. [^13] Fowlie et.al. [9] estimate a price elasticity of demand for electricity in the U.S. of -2.06 for with a range of alternative elasticities around -1 (see Table 2). These results show that an increase in electricity demand will be split between increases in both price and quantity.
2.2 Externalities costs that decrease with hashrate
@@
We review three vectors of attack on a PoW blockchain that decrease in profitability as hashrate increases.
@@
Double-spend attack A double-spend attack requires the attacker to build an alternative chain that overtakes the incumbent canonical chain at in terms of cumulative mining puzzle difficulty. An attacker with over 50% of total hashrate will overtake the incumbent canonical chain at some block with probability measure 1. At less than 50% share, the attacker’s probability of success is less than half. [^14] See Gervais et.al. [11] for simulations of attacks using different ratios of attacker hashrate to miner hashrate. The frequency of success is increasing in attacker hashrate and the average number of blocks required to overtake the incumbent canonical chain is decreasing in attacker hashrate. The amount of attacker hashrate required to successfully fork the blockchain, and its cost, is increasing in the level of honest miner hashrate. Budish (2018) [3] and Aronoff et.al. [1] point out that the cost of a successful attack is at least partially offset by the block rewards that accrue to the attacker. The possibility that a victim of a double-spend attack retaliates by mining on the incumbent canonical chain leads to a war of attrition between the attacker and the victim (Moroz et.al. [16]). Aronoff et.al. [1]) point out that the outcome of a war of attrition depends on equilibrium refinements and exigent circumstances, even when the costs between players differ. Nevertheless, an increase in the cost of launching the attack will act as a deterrent by increasing the sunk cost that the attacker loses if it is unsuccessful. Equation 1 is the attacker’s un-discounted ex-ante expected profit when block rewards equal hashrate cost. The key observation is that, for any probability of success, the attacker’s expected profit is declining in honest miner hashrate.
@@
Ex-ante Expected $\text{Ex-ante Expected Profit}\;=\;Prob(success)V-\\ \{1-Prob(success)\}\text{E[hashrate cost per block]E[number of blocks]}$ (1)
@@
2.3 The target hashrate interval
Figure 2 is a conceptual depiction of the curves representing network security and energy externalities cost. The horizontal axis is the hashrate of miners in the Bitcoin network. The vertical axis represents the cost of security and energy externalities. Total cost is the summation of these curves at each level of hashrate. The goal of Targeted is minimize the total cost of security and energy externalities. The negative slope of the network security curve reflects that the network becomes less vulnerable as hashrate increases. Network security vanishes as hashrate approaches zero, since the cost of acquiring 51% hashrate is negligible (equivalent to unbounded security cost). The cost of attacking the network becomes extremely expensive at very high levels (equivalent to negligible security cost). A mild smoothness assumption implies that the network security curve is convex in hashrate. I assume the unit hashrate cost is either constant or increases with electricity consumption (and hashrate). The shape of the social cost of carbon emissions is convex in electricity consumption (and hashrate).161616 ”The best evidence is that impacts of climate change will be non-linear and cumulative.” (William Nordhaus Nobel Prize lecture in economics [18]). Deitz et.al. [7] identify the ”three essential elements of the climate problem – the endogeneity of growth, the convexity of damage and climate risk”. The energy externality cost is the summation of these two costs, which is convex. The sum of the two curves is convex, which assures the existence of an interval in which total cost is minimized. A reduction in per hash enery cost induced by improvements in ASIC efficiency and/or electricity cost will shift the network security curve to the right. A reduction in per hash carbon emissions from an increase in the use of clean energy will shift the hashrate/energy externalities cost curve to the right.
Figure 2: The Target Hashrate Interval

### v4

# Targeted Nakamoto: A Bitcoin Protocol to Balance Network Security and Carbon Emissions

@@ Abstract
In a Proof-of-Work blockchain like Bitcoin, mining hashrate increases with block rewards. Higher hashrate reduces vulnerability to attacks (lowering network security cost) but raises carbon emissions and electricity use (raising environmental cost), reflecting a tradeoff with a hashrate point or interval that minimizes total cost. Targeted Nakamoto is a protocol that incentivizes miners to home hashrate in on the minimum cost range by capping block rewards when above target and imposing a block reward floor when below. Monetary neutrality is maintained by proportional adjustments in spending potential among UTXO holders, offsetting the additions and subtractions to the block reward caused by the protocol. Bitcoin faces two conflicting existential risks: currently, high mining energy consumption invites political backlash; in the future, reductions in miner rewards will cause hashrate to decline, lowering the cost of an attack. Targeted Nakamoto balances these concerns by guiding hashrate towards a chosen target that places an effective floor on the cost of attack and a ceiling on carbon emissions.
@@ 1 Introduction
Proof-of -Work (PoW) blockchain cryptocurrencies like Bitcoin require the application of computing power by miners to operate the network. Miners assemble blocks and compete to solve a puzzle set by the code. The number of puzzle guesses (one hash per guess) a mining computer makes in a specified interval of time is its hashrate, which consumes electricity. Miners pay the direct cost of electricity. There are, in addition, two externalities created by mining activity. One externality is the vulnerability of the blockchain network to attack. The degree of vulnerability is related to the cost of forking the blockchain, which in turn is related to the cost of hashrate and the volume of hashrate employed by miners. The higher is hashrate, the higher is the cost of forking a PoW blockchain, which implies that network security is increasing in hashrate (equivalently, security cost is decreasing in hashrate). The other externality arises from the electricity consumed by hashrate. Miner demand for electricity increases both the price and aggregate consumption of electricity. The former is a cost imposed on competing users of electricity. The latter generates carbon emissions, which has a social cost. These electricity related costs are increasing in hashrate. The interaction of security risk and carbon emissions imply there is a tradeoff in total cost, defined as the sum of the two externalities costs, at different levels of hashrate. An increase in hashrate reduces the vulnerability of the network (equivalently, security cost is decreased), while it increases carbon emissions and creates upward pressure on the price of electricity. There may be a point or interval of hashrate in which total cost is minimized. Current PoW protocols, which I refer to as ”Nakamoto”, do not restrict hashrate.111 Our focus in this paper is Bitcoin, but the protocol applies to any PoW cryptocurrency.
The key lever upon which the new protocol is built is the incentive effect of the block reward. An increase in the USD value of the block reward creates an incentive for miners to apply more hashrate (assuming the dollar is the currency in which miners calculate profit). A decrease in the USD value of the block reward has the opposite effect. Equilibrium hashrate can range from zero to unbounded. Figure 5 shows the tight correlation over time between percentage change in the USD value of the Bitcoin block reward and the percentage change in energy consumption from Bitcoin mining.
@@
Figure 1: Bitcoin Energy Consumption vs USD Mining Rewards. Source: CBECI44footnotemark: 4and Coinmetrics. [^5] https://ccaf.io/cbnsi/cbeci
55footnotetext: https://coinmetrics.io
@@
Targeted Nakamoto (hereafter ”Targeted”) modifies Nakamoto to create an incentive for miners to push hashrate toward an interval whenever it strays outside the interval. The miner’s minimum block reward (in BTC) is increased when hashrate is below the interval, which induces miners to apply more hashrate. The miner’s maximum block reward (in BTC) is decreased when hashrate is above the interval, which induces miners to apply less hashrate. Monetary neutrality is maintained by offsetting the increase and decrease in the miner’s block reward by opposite adjustments to the spending potential of addresses with UTXO value, in proportion to the value held by each address.
1.1 Balancing Competing Risks
Targeted balances the ecological threat from electricity consumed by hashrate against the threat to network security from low hashrate. The former is a subject of current political controversy, while the latter will materialize in the future when the USD value of the block reward declines as a result of the halving of the coinbase.  [^6] Historically, the trend USD/BTC exchange rate has risen sufficiently to offset the halving of the coinbase. However, this cannot continue indefinitely for two reasons. First, coinbase will eventually be zero. Second, maintaining a fixed USD value of the coinbase paid to miners as halving continues would require the inflation adjusted USD to BTC exchange rate to double every four years. This implies a growth rate that exceeds US (and world) GDP by orders of magnitude, which is dynamically unsustainable (Tirole [21]). Transactors will have little incentive to compensate by increasing fees, which are paid to obtain a priority in the qeue of mempool transactions. The decline in block reward will elicit a decline in hashrate (Equation 4). Targeted compensates by increasing the block reward, and does it without increasing overall UTXO spending potential.
1.2 Links to code and web API (by Kristian Praizner)
@@
Attached are links to the code-base for controlling Hashrate and a Web API with an empirically estimated hashrate function and a module that enables the user to choose a target hashrate interval and control parameters and to run counterfactual experiments with alternative time-paths of model variables.
@@ 1.4 The research question and design constraints
Targeted augments Nakamoto by creating an incentive for miners to home in on a hashrate interval. The augmentation comes close to adhering to formal constraints which reflect core values held by the Bitcoin community. [^9] What constitutes a core feature is subjective. I developed the list herein after consultation with a number of participants in the Bitcoin network, including developers, miners and exchange operators. The constraints are partitioned into the consensus protocol layer (C constraints) and the transaction protocol layer (T- constraints).
@@ 1.5 Roadmap
The rest of the paper is organized as follows. Section 2 provides an overview of the two types of cost, the sum of which Targeted is designed to minimize; environmental costs that increase with hashrate and network sercurity costs that decrease with hashrate. Section 3 casts Nakamoto and Targeted in a mechanism design framework in order to articulate the key features of each protocol and their interaction and to show how Targeted augments Nakamoto. Section 4 shows how the ratio of mining puzzle difficulty to blockchain growth rate is a sufficient statistic for hashrate and then derives the expression for equilibrium hashrate as a function of, inter alia, the value of the block reward. Section 5 uses the relationship between the puzzle difficulty/growth rate ratio and hashrate, along with hashrate equilibrium to construct the policy for adjusting the miner’s block reward to incentivize miners to push hashrate toward the target hashrate interval. Section 6 demonstrates that the block reward policy is consistent with Nash equilibrium at the targeted hashrate and demonstrates that the hashrate control algorithm dynamically pushes hashrate toward the target interval. Section 7 states a protocol that renders the aggregate money supply and the distribution of spending potential unaffected by the Hashrate Control Algorithm. The paper concludes with a summary of the key properties of Targeted.
2 Hashrate Externalities
There are intensive and extensive margins related to hashrate in Bitcoin mining. The intensive margin is the equilibrium hashrate cost paid by miners. The extensive margins are the social cost of $CO^{2}$ emissions from the electricity used to generate hashrate and the network security risks arising from the cost of an attack that forks the blockchain. The former is increasing and the latter is decreasing in aggregate miner hashrate. The causal link from hashrate to these variables is determined by ASIC efficiency and electricity cost. Equation 1 displays the relationship.
$\underset{\text{per-hash cost}}{\underbrace{cost/N}}=\underset{\text{electricity cost}}{\underbrace{(cost/KwH)}}\div\underset{\text{ASIC efficiency}}{\underbrace{(N/KwH)}}$ (1)
Where $N$ is hashrate (defined in Section 4), $KwH$ is kilowatt hours. Environmental cost is an increasing function of hashrate as follows.
$\text{environmental cost}\underset{\text{external function}}{\underbrace{\longleftarrow}}CO^{2}\underset{\text{physical law}}{\underbrace{\longleftarrow}}KwH=\textbf{{N}}\div\text{ASIC efficiency}$
Network security cost is a decreasing function of hashrate as follows.
$\text{network security cost}\underset{\text{external function}}{\underbrace{\longleftarrow}}\text{mining cost}=\textbf{{N}}\times\text{electricity cost}\div\text{ASIC efficiency}$
Figure 2 displays the relationship between environmental cost and network security cost on the vertical axes, and hashrate on the horizontal axis. is a upward sloped curve and network security is a downward sloped curve. We define optimal hashrate interval as the point, or interval, where the sum of the costs are minimized. The goal of Targeted is to incentivize miners to home in on this interval.
Figure 2: Target Hashrate Interval
2.1 Dependencies
Assuming the external functions are stable over time, environmental risk and network security are conditional on ASIC efficiency and electricity cost. [^10] The function connecting electricity to $CO^{2}$ is a physical law. The function connecting forking cost to attack incentives is subject to variables that lie outside our analysis, such as, inter alia, interest rates and the legal environment. An increase in ASIC efficiency will shift both curves rightward and increase the optimal hashrate target. An increase in electricity cost will shift the security curve rightward while leaving the environmental risk curve unaffected, and reduce the optimal hashrate. ASIC efficiency and electricity cost can be estimated and oracles can be implemented to adjust the hashrate target. Additionally, there is a causal channel running from hashrate to electricity cost, whereby the increased demand for electricity places upward pressure on the price. Fowlie et.al. [9] estimate a price elasticity of demand for electricity in the U.S. of -2.06 with a range of alternative elasticities around -1, which implies that an increase in electricity demand will be split between increases in both price and quantity.
Less clear is how to measure the external functions. There is a literature on measuring the social cost of $CO^{2}$ emissions. Nordhaus [18]), Deitz et.al. and Stern [7] are important contributions and Chen et.al. [14] model the flow of carbon emissions from multiple energy sources. Nevertheless, the shape and magnitude of the relationship between $CO^{2}$ emissions and social cost is a matter of ongoing debate. There has not been any substantive work, to our knowledge, connecting mining cost to quantifiable network security. [^11] Appendix 2.2 discusses the relationship between mining cost and network attack vectors. A market in Bitcoin security risk, with payoffs tied to attack events, could resolve the matter. But for now, this function can only be guessed at. The practical implementation of Targeted will require the formation of a consensus among Bitcoin network participants on the external functions. This may require negotiations that will involve compromise over differing opinions and economic interests.
2.2 Network Security
@@
To elucidate the relationship between hashrate cost and network security, we review three vectors of attack on a PoW blockchain that decrease in profitability as hashrate cost increases.
@@
Double-spend attack A double-spend attack requires the attacker to build an alternative chain that overtakes the incumbent canonical chain at in terms of cumulative mining puzzle difficulty. An attacker with over 50% of total hashrate will overtake the incumbent canonical chain at some block with probability measure 1. At less than 50% share, the attacker’s probability of success is less than half. [^12] See Gervais et.al. [11] for simulations of attacks using different ratios of attacker hashrate to miner hashrate. The frequency of success is increasing in attacker hashrate and the average number of blocks required to overtake the incumbent canonical chain is decreasing in attacker hashrate. The amount of attacker hashrate required to successfully fork the blockchain, and its cost, is increasing in the level of honest miner hashrate. Budish [3] and Aronoff et.al. [1] point out that the cost of a successful attack is at least partially offset by the block rewards that accrue to the attacker. The possibility that a victim of a double-spend attack retaliates by mining on the incumbent canonical chain leads to a war of attrition between the attacker and the victim (Moroz et.al. [16]). Aronoff et.al. [1] show that the outcome of a war of attrition depends on equilibrium refinements and exigent circumstances, even when the costs between players differ. Nevertheless, an increase in the cost of launching the attack will act as a deterrent by increasing the sunk cost that the attacker loses if it is unsuccessful. Equation 2 is the attacker’s un-discountedex-ante expected profit when block rewards equal hashrate cost. The key observation is that, for any probability of success, the attacker’s expected profit is declining in honest miner hashrate.
@@
$\text{Ex-ante Expected Profit}\;=\;Prob(success)V-\\ \{1-Prob(success)\}\text{E[hashrate cost per block]E[number of blocks]}$ (2)
@@
Finally, Carlesten et.al. [4] argued that as the coinbase approaches zero and miners become dependent on transaction fees, the decline and variability of the fees will incentivize selfish mining, which could disrupt the liveness of the blockchain and possibly push down hashrate (and thereby mining cost).